# Revisiting the Reliability of Language Models in Instruction-Following

### v1

@@ A.1 Benchmarking Instruction-Following
SysBench (qin2024sysbench-system-message-following-benchmark-with-constraints). This work echoes the real-world user practices involving system prompts. They evaluate the instruction-following ability of LLMs, involving 500 system prompts requesting six types of constraints.
@@
MMMT-IF (epstein2024mmmt-if-multimodal-multi-turn-instruction-following). This work extends the instruction-following evaluation to multi-modal and multi-turn settings. The MMMT-IF benchmark contains 71 test cases, with image inputs and multiple constraints scattered across chat.
@@
RealInstruct (ferraz2024llm-realinstruct-real-world-instruction-following-emnlp). This paper introduces real-world user queries as a source for instruction-following test cases. The authors meticulously curate instruction candidates from ShareGPT and then decompose them into tasks, contexts, and constraints. The final test set includes 302 instances, each of which typically contains multiple constraints.
@@
Multi-IF (he2024multi-if-multi-turn-and-multilingual-instructions-following-meta). This work extends instruction-following evaluation to multi-turn and multilingual settings. The Multi-IF benchmark is composed of 4,501 multilingual conversations, constructed by augmenting IFEval with multi-turn interactions and translations.
@@
LIFBench (wu-etal-2025-lifbench-instruction-following-in-long-context). This work studies instruction-following abilities of LLMs in long-context tasks, e.g., processing multiple documents. The benchmark consists of 2,766 instructions, involving long-context tasks and 6 types of instruction-following capabilities.
@@
VFF (wang2025verifiable-vff-naacl-format-control). This work exclusively explores the format-constraint following abilities of LLMs. It constructs 21K test cases with varying levels of difficulty by combining one to three verifiable constraints per instruction. The constraint types include limited word count, limited content, limited punctuation, limited structure, limited grammar, and specific number formatting.
@@
StructFlowBench (li2025structflowbench-structured-flow-benchmark-for-multi-turn-instruction-following). This work explores the instruction-following performance of LLMs in multi-turn scenarios, emphasizing the structural dependencies that arise across turns. The benchmark incorporates eight types of constraints and six categories of structural flows: follow-up, refinement, recall, expansion, summary, and unrelatedness. In total, it includes 155 dialogues, evaluated via LLM-as-a-Judge.
@@
PBIF zeng2025order-position-bias-in-multi-constraint-instruction-following. This work investigates how the ordering of multiple constraints influences the instruction-following performance of LLMs. It identifies an intriguing phenomenon: LLMs often perform better when the constraints are presented in a hard-to-easy order.
@@
CodeIF (yan2025codeif). This work introduces a benchmark for evaluating LLMs’ instruction-following in code generation, covering tasks like synthesis, debugging, refactoring, and explanation. It comprises 1,200 test cases across 8 categories and 50 sub-instructions, spanning multiple programming languages.
@@
WildIFEval (lior2025wildifeval). This work curates a set of 12K complex instructions collected from real-world user queries on Chatbot Arena (chiang2024chatbot-arena-an-open-platform-for-evaluating-llms-by-human-preference-icml). Each instruction is then decomposed into 8 categories of constraints, and LLM-as-a-Judge is employed for automated evaluation of instruction-following.
@@
XIFBench (li2025xifbench-multilingual-instruction-following). This work primarily explores the instruction-following abilities of LLMs across different language settings, from high-resource (English and Chinese) to low-resource (Hindi). It starts with five categories of constraints: content, style, situation, format, and numerical. By composing between one and five constraints into each seed instruction, the authors construct a total of 465 hard instructions. These are further extended into multiple languages through translation, resulting in a multilingual benchmark.
@@
CodeIF-Bench (wang2025codeif-bench-interactive-code-generation). This work focuses on multi-turn interactive code generation. It covers 9 types of verifiable instructions aligned with real-world software development requirements for test-case validation, and supports both Static and Dynamic Conversation settings.
@@
MulDimIF ye2025multi-dimensional-constraint-framework-fudan. This work proposes a multi-dimensional constraint framework, involving three types of constraint formatting patterns (Example, Listing, and Incorporation), four categories of constraints, and four difficulty levels. In total, this benchmark consists of 1,200 code-verifiable instructions.
@@
MathIF (fu2025scalingreasoninglosingcontrol) is a dedicated benchmark for evaluating instruction-following in mathematical reasoning tasks. It contains 420 evaluation samples derived from 15 Python-verifiable constraints across 4 categories, applied to math problems of varying difficulty.
@@
AGENTIF (qi2025agentifbenchmarkinginstructionfollowing) evaluates LLMs’ instruction-following abilities in agentic scenarios with long, complex, and constraint-heavy instructions. It comprises 707 human-annotated instructions drawn from 50 real-world agentic tasks.
@@
LIFEBench (zhang2025lifebench-evaluating-length-instruction). This work is conducted with a focus on length constraints, ranging from 16 to 8192 words. It contains 10,800 instructions, in English or Chinese, requesting responses with length satisfying three types of relations (At Most, At Least, and Equal To).
@@
MaXIFE (liu2025maxifemultilingualcrosslingualinstruction). This work focuses on multilingual and cross-lingual scenarios. It consists of parallel data in 23 languages, with each record combining a Basic Question and 1–3 Instructions, totaling 18,285 entries.
@@
Ordered CommonGen (song2025ifircomprehensivebenchmarkevaluating) introduces a benchmark that evaluates LLMs’ ability to generate concepts in a specified order. It contains 27,648 instances from 192 concept sets and 6 instruction templates, using ordered coverage to measure adherence to the specified sequence.
IFBench (pyatkin2025generalizing-if-bench). This work highlights the limited generalization of LLMs’ instruction-following abilities across constraint types, thus falling into comprehensiveness-oriented reliability. They curate a new set of 300 instructions, involving 58 new constraint types beyond the 25 types in IFEval.
IFScale (jaroslawicz2025instructionsllmsfollowonce). This work introduces a benchmark for evaluating instruction-following performance as instruction density increases in professional business report generation. Each task contains 10–500 keyword-inclusion instructions drawn from a curated 500-term business vocabulary, with density increasing in steps of 10.
IFEvalCode (yang2025ifevalcodecontrolledcodegeneration) introduces a multilingual benchmark for controlled code generation, evaluating LLMs’ ability to follow instructions beyond correctness in coding scenarios. It includes 1,600 samples in 8 languages with paired Chinese–English queries, assessing both correctness and instruction adherence.
@@
    CoDI-Eval (chen2024benchmarking-CoDI-Eval-aaai-diversified-instructions): This benchmark mainly focuses on the controllability of LLMs in generation tasks. This work advocates an instruction diversification process to synthesize diverse forms of constraint expression (i.e., the rewriting step). They aim to achieve finer-grained coverage of instructions rather than to study the reliability issue. What’s more, they fail to account for inter-prompt relationships.
@@
    SIFo (chen2024sifo-sequential-instruction-following) & PBIF (zeng2025order-position-bias-in-multi-constraint-instruction-following): These two works investigate the position bias problem of LLMs in the context of instruction-following. They examine how LLMs’ compliance changes when the order of constraints within prompts is altered. This phenomenon can be viewed as a special case of our rephrasing-based augmentation approach.
@@
    LIFEBench (zhang2025lifebench-evaluating-length-instruction): This work specifically studies length constraints. Their study is analogous to our pilot study in Section˜2.1. However, they focus on how long-context LLMs follow length constraints rather than on the general reliability.
@@
    IFScale (jaroslawicz2025instructionsllmsfollowonce): This work studies how many instructions LLMs can simultaneously follow. For the research question, the study also relates to reliability, but it stress-tests LLMs along the complexity-oriented dimension.
@@ A.2 Evaluating LLMs’ Reliability
There exist works revealing LLMs’ sensitivity to prompt wordings (sclar2024quantifying-how-i-learned-to-start-worrying-about-prompt-formatting-classification-tasks). cao2024worst-prompt-performance-of-llms specifically study the worst performance across semantically equivalent case-level queries. mizrahi2024state-multi-prompt-llm-evaluation advocate for multi-prompt LLM evaluation, which shares the same methodology as our curation of cousin prompts. Based on these works, we move beyond studying simply prompt sensitivity issues (the rephrasing augmentation may reflect) and towards more general nuance-oriented reliability.
A critical distinction must be drawn between the reliable@k metric and sampling responses for one prompt independently kk times. yao2025taubench-tool-calling-agent introduce a new metric, passˆk, for evaluating agentic scenarios. While passˆk measures reliability across kk independent trials using the same prompt (often with temperature >0>0), our reliable@k metric evaluates consistency across distinct cousin prompts with nuanced variations. To quantitatively compare these two dimensions, we conducted an experiment sampling each IFEval prompt 10 times at temperature 1 to compute passˆ10, contrasting it with our reliable@k on IFEval++.
@@ Appendix C Model Details
In this work, we include a total of 36 large language models to establish a representative benchmark. Our selection aims to comprehensively capture the development trends of LLMs in terms of reliability. To this end, we include models that are widely recognized for their performance or influence within the research and developer communities, covering a diverse range of providers, architectures, and scales. For those proprietary models, we access their LLM services via API, specifically the OpenAI-compatible interface. For the open-source ones, we locally deploy them using vLLM (kwon2023vllm) for inference. We use the instruction-tuned or chat version and adopt official chat templates. See Table˜8 for the detailed information.
@@ D.1 Impact of the System Prompt
Meanwhile, we notice that it is a common practice for service providers or end users to employ system prompts tailored for personalized LLM services (ge2024scaling-persona-hub; qin2024sysbench-system-message-following-benchmark-with-constraints). We are also curious about the impact of system prompt setting on the models’ reliability in instruction-following. We include six settings as listed in Table˜9: the default, three persons sampled from ge2024scaling-persona-hub, python coder, and encouraging better instruction-following. Except for the system prompt setting, other decoding strategies remain consistent.
@@ D.3 Pilot Experiments on Qwen3-32B
We extend our pilot experiments to Qwen3-32B. As shown in Figure˜8, the instability persists in the larger-scale LLM. Overall, for the more challenging Around relation, the LLM is far less reliable, consistent with the findings of zhang2025lifebench-evaluating-length-instruction. The instability reveals the vulnerability of LLMs in nuance-oriented reliability, motivating us to study this critical yet underexplored dimension.
@@ E.1 Training Details
We include two dataset settings in the SFT experiments: (1) general-domain instruction-following, for which we use the representative alpaca (alpaca) as the data source; (2) self-curated datasets tailored for improving reliability, for which we employ another pool of cousin prompts for the IFEval. Across the two settings, the system prompt is consistently set to the default one for both training and evaluation.
@@
Additionally, we advance an exploratory attempt around distinct training methods. We include the direct preference optimization (DPO) algorithm (rafailov2024dpo), an alignment method that directly optimizes a model’s parameters toward preferred responses without requiring explicit reward modeling. Unlike reinforcement learning from human feedback (RLHF) (ouyang2022instructgpt), which relies on a separately trained reward model and policy optimization, DPO leverages pairwise preference data. This encourages the model to assign higher relative likelihood to preferred responses compared to rejected ones, enabling alignment without explicit reward modeling. The experiments are conducted on the same Cousin Prompts dataset. For DPO, the rejected samples are collected from LLaMA-3.3-70B-Instruct. When multiple rejected responses are available, we randomly select from those that fail to meet the evaluation criteria. If none explicitly fail, we instead choose the response with the largest character count, as it intuitively represents the least efficient way of satisfying the requirements. All other training configurations are shared across the two training methods, including the learning rate.
@@
We leave this non-trivial exploration of training methods in the Appendix, as we cannot draw conclusions about best practices. Future works may extend this investigation to other training paradigms, e.g., reinforcement learning from verifiable rewards (lambert2024tulu-3-rlvr; guo2025deepseek-r1; zou2025eifbench).

### v2

@@ A.1 Benchmarking Instruction-Following
SysBench (Qin et al., 2025). This work echoes the real-world user practices involving system prompts. They evaluate the instruction-following ability of LLMs, involving 500 system prompts requesting six types of constraints.
@@
MMMT-IF (Epstein et al., 2024). This work extends the instruction-following evaluation to multi-modal and multi-turn settings. The MMMT-IF benchmark contains 71 test cases, with image inputs and multiple constraints scattered across chat.
@@
RealInstruct (Ferraz et al., 2024). This paper introduces real-world user queries as a source for instruction-following test cases. The authors meticulously curate instruction candidates from ShareGPT and then decompose them into tasks, contexts, and constraints. The final test set includes 302 instances, each of which typically contains multiple constraints.
@@
Multi-IF (He et al., 2024b). This work extends instruction-following evaluation to multi-turn and multilingual settings. The Multi-IF benchmark is composed of 4,501 multilingual conversations, constructed by augmenting IFEval with multi-turn interactions and translations.
@@
LIFBench (Wu et al., 2025). This work studies instruction-following abilities of LLMs in long-context tasks, e.g., processing multiple documents. The benchmark consists of 2,766 instructions, involving long-context tasks and 6 types of instruction-following capabilities.
@@
VFF (Wang et al., 2025b). This work exclusively explores the format-constraint following abilities of LLMs. It constructs 21K test cases with varying levels of difficulty by combining one to three verifiable constraints per instruction. The constraint types include limited word count, limited content, limited punctuation, limited structure, limited grammar, and specific number formatting.
@@
StructFlowBench (Li et al., 2025a). This work explores the instruction-following performance of LLMs in multi-turn scenarios, emphasizing the structural dependencies that arise across turns. The benchmark incorporates eight types of constraints and six categories of structural flows: follow-up, refinement, recall, expansion, summary, and unrelatedness. In total, it includes 155 dialogues, evaluated via LLM-as-a-Judge.
@@
PBIF Zeng et al. (2025). This work investigates how the ordering of multiple constraints influences the instruction-following performance of LLMs. It identifies an intriguing phenomenon: LLMs often perform better when the constraints are presented in a hard-to-easy order.
@@
CodeIF (Yan et al., 2025). This work introduces a benchmark for evaluating LLMs’ instruction-following in code generation, covering tasks like synthesis, debugging, refactoring, and explanation. It comprises 1,200 test cases across 8 categories and 50 sub-instructions, spanning multiple programming languages.
@@
WildIFEval (Lior et al., 2025). This work curates a set of 12K complex instructions collected from real-world user queries on Chatbot Arena (Chiang et al., 2024). Each instruction is then decomposed into 8 categories of constraints, and LLM-as-a-Judge is employed for automated evaluation of instruction-following.
@@
XIFBench (Li et al., 2025b). This work primarily explores the instruction-following abilities of LLMs across different language settings, from high-resource (English and Chinese) to low-resource (Hindi). It starts with five categories of constraints: content, style, situation, format, and numerical. By composing between one and five constraints into each seed instruction, the authors construct a total of 465 hard instructions. These are further extended into multiple languages through translation, resulting in a multilingual benchmark.
@@
CodeIF-Bench (Wang et al., 2025a). This work focuses on multi-turn interactive code generation. It covers 9 types of verifiable instructions aligned with real-world software development requirements for test-case validation, and supports both Static and Dynamic Conversation settings.
@@
MulDimIF Ye et al. (2025). This work proposes a multi-dimensional constraint framework, involving three types of constraint formatting patterns (Example, Listing, and Incorporation), four categories of constraints, and four difficulty levels. In total, this benchmark consists of 1,200 code-verifiable instructions.
@@
MathIF (Fu et al., 2025) is a dedicated benchmark for evaluating instruction-following in mathematical reasoning tasks. It contains 420 evaluation samples derived from 15 Python-verifiable constraints across 4 categories, applied to math problems of varying difficulty.
@@
AGENTIF (Qi et al., 2025) evaluates LLMs’ instruction-following abilities in agentic scenarios with long, complex, and constraint-heavy instructions. It comprises 707 human-annotated instructions drawn from 50 real-world agentic tasks.
@@
LIFEBench (Zhang et al., 2025b). This work is conducted with a focus on length constraints, ranging from 16 to 8192 words. It contains 10,800 instructions, in English or Chinese, requesting responses with length satisfying three types of relations (At Most, At Least, and Equal To).
@@
MaXIFE (Liu et al., 2025). This work focuses on multilingual and cross-lingual scenarios. It consists of parallel data in 23 languages, with each record combining a Basic Question and 1–3 Instructions, totaling 18,285 entries.
@@
Ordered CommonGen (Song et al., 2025) introduces a benchmark that evaluates LLMs’ ability to generate concepts in a specified order. It contains 27,648 instances from 192 concept sets and 6 instruction templates, using ordered coverage to measure adherence to the specified sequence.
IFBench (Pyatkin et al., 2025). This work highlights the limited generalization of LLMs’ instruction-following abilities across constraint types, thus falling into comprehensiveness-oriented reliability. They curate a new set of 300 instructions, involving 58 new constraint types beyond the 25 types in IFEval.
IFScale (Jaroslawicz et al., 2025). This work introduces a benchmark for evaluating instruction-following performance as instruction density increases in professional business report generation. Each task contains 10–500 keyword-inclusion instructions drawn from a curated 500-term business vocabulary, with density increasing in steps of 10.
IFEvalCode (Yang et al., 2025b) introduces a multilingual benchmark for controlled code generation, evaluating LLMs’ ability to follow instructions beyond correctness in coding scenarios. It includes 1,600 samples in 8 languages with paired Chinese–English queries, assessing both correctness and instruction adherence.
@@
    CoDI-Eval (Chen et al., 2024c): This benchmark mainly focuses on the controllability of LLMs in generation tasks. This work advocates an instruction diversification process to synthesize diverse forms of constraint expression (i.e., the rewriting step). They aim to achieve finer-grained coverage of instructions rather than to study the reliability issue. What’s more, they fail to account for inter-prompt relationships.
@@
    SIFo (Chen et al., 2024b) & PBIF (Zeng et al., 2025): These two works investigate the position bias problem of LLMs in the context of instruction-following. They examine how LLMs’ compliance changes when the order of constraints within prompts is altered. This phenomenon can be viewed as a special case of our rephrasing-based augmentation approach.
@@
    LIFEBench (Zhang et al., 2025b): This work specifically studies length constraints. Their study is analogous to our pilot study in Section˜2.1. However, they focus on how long-context LLMs follow length constraints rather than on the general reliability.
@@
    IFScale (Jaroslawicz et al., 2025): This work studies how many instructions LLMs can simultaneously follow. For the research question, the study also relates to reliability, but it stress-tests LLMs along the complexity-oriented dimension.
@@ A.2 Evaluating LLMs’ Reliability
There exist works revealing LLMs’ sensitivity to prompt wordings (Sclar et al., 2024). Cao et al. (2024) specifically study the worst performance across semantically equivalent case-level queries. Mizrahi et al. (2024) advocate for multi-prompt LLM evaluation, which shares the same methodology as our curation of cousin prompts. Based on these works, we move beyond studying simply prompt sensitivity issues (the rephrasing augmentation may reflect) and towards more general nuance-oriented reliability.
A critical distinction must be drawn between the reliable@k metric and sampling responses for one prompt independently kk times. Yao et al. (2025) introduce a new metric, passˆk, for evaluating agentic scenarios. While passˆk measures reliability across kk independent trials using the same prompt (often with temperature >0>0), our reliable@k metric evaluates consistency across distinct cousin prompts with nuanced variations. To quantitatively compare these two dimensions, we conducted an experiment sampling each IFEval prompt 10 times at temperature 1 to compute passˆ10, contrasting it with our reliable@k on IFEval++.
@@ Appendix C Model Details
In this work, we include a total of 36 large language models to establish a representative benchmark. Our selection aims to comprehensively capture the development trends of LLMs in terms of reliability. To this end, we include models that are widely recognized for their performance or influence within the research and developer communities, covering a diverse range of providers, architectures, and scales. For those proprietary models, we access their LLM services via API, specifically the OpenAI-compatible interface. For the open-source ones, we locally deploy them using vLLM (Kwon et al., 2023) for inference. We use the instruction-tuned or chat version and adopt official chat templates. See Table˜8 for the detailed information.
@@ D.1 Impact of the System Prompt
Meanwhile, we notice that it is a common practice for service providers or end users to employ system prompts tailored for personalized LLM services (Ge et al., 2024; Qin et al., 2025). We are also curious about the impact of system prompt setting on the models’ reliability in instruction-following. We include six settings as listed in Table˜9: the default, three persons sampled from Ge et al. (2024), python coder, and encouraging better instruction-following. Except for the system prompt setting, other decoding strategies remain consistent.
@@ D.3 Pilot Experiments on Qwen3-32B
We extend our pilot experiments to Qwen3-32B. As shown in Figure˜8, the instability persists in the larger-scale LLM. Overall, for the more challenging Around relation, the LLM is far less reliable, consistent with the findings of Zhang et al. (2025b). The instability reveals the vulnerability of LLMs in nuance-oriented reliability, motivating us to study this critical yet underexplored dimension.
@@ E.1 Training Details
We include two dataset settings in the SFT experiments: (1) general-domain instruction-following, for which we use the representative alpaca (Taori et al., 2023) as the data source; (2) self-curated datasets tailored for improving reliability, for which we employ another pool of cousin prompts for the IFEval. Across the two settings, the system prompt is consistently set to the default one for both training and evaluation.
@@
Additionally, we advance an exploratory attempt around distinct training methods. We include the direct preference optimization (DPO) algorithm (Rafailov et al., 2024), an alignment method that directly optimizes a model’s parameters toward preferred responses without requiring explicit reward modeling. Unlike reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022), which relies on a separately trained reward model and policy optimization, DPO leverages pairwise preference data. This encourages the model to assign higher relative likelihood to preferred responses compared to rejected ones, enabling alignment without explicit reward modeling. The experiments are conducted on the same Cousin Prompts dataset. For DPO, the rejected samples are collected from LLaMA-3.3-70B-Instruct. When multiple rejected responses are available, we randomly select from those that fail to meet the evaluation criteria. If none explicitly fail, we instead choose the response with the largest character count, as it intuitively represents the least efficient way of satisfying the requirements. All other training configurations are shared across the two training methods, including the learning rate.
@@
We leave this non-trivial exploration of training methods in the Appendix, as we cannot draw conclusions about best practices. Future works may extend this investigation to other training paradigms, e.g., reinforcement learning from verifiable rewards (Lambert et al., 2024; Guo et al., 2025; Zou et al., 2025).